OpenStudy (anonymous):

Medal A volleyball playoff competition has eliminations each round. The table below shows the number of teams in each round of the tournament: Round (x) 1 2 3 4 Teams f(x) 16 8 4 2 Compute the average rate of change of f(x) from x = 1 to x = 4 and identify the meaning of that rate. negative three over fourteen; on average, there were 3 fewer teams every 14 rounds negative fourteen over three; on average, there were 14 fewer teams every 3 rounds 3; there were 3 rounds between rounds 1 and 4 14; there were 14 fewer team between rounds 1 and 4

OpenStudy (solomonzelman):

\(\large\color{blue}{ \huge{ \begin{array}{| l | c | r |} \hline \scr~~~x~~ & \scr y \\ \hline \scr~~~1~ & \scr 16 \\ \hline \scr~~~2~ & \scr 8 \\ \hline \scr~~~3~ & \scr 4 \\ \hline \scr~~~4~ & \scr 2 \\ \hline \end{array} } }\) this is your chart ?

OpenStudy (anonymous):

yes

OpenStudy (solomonzelman):

when x=1, y=16 when x=4, y=2 (this is simply based on the chart) So, your to points are (1,16) and (4,2) are your two points. \(-------------------------\)

OpenStudy (solomonzelman):

THE SLOPE FORMULA: \(\LARGE \color{black}{ \displaystyle {\rm m}=\frac{\color{blue}{{\rm y}_1}-\color{red}{{\rm y}_2}}{\color{green}{{\rm x}_1}-\color{darkgoldenrod}{{\rm x}_2}} }\) where \(\LARGE \color{black}{ \displaystyle {\rm m} }\) is the slope \(\Large\color{black}{ \displaystyle (\color{green}{{\rm x}_1}~,~~\color{blue}{{\rm y}_1}) }\) and \(\Large\color{black}{ \displaystyle (\color{darkgoldenrod}{{\rm x}_2}~,~~\color{red}{{\rm y}_2}) }\) are your two points.

OpenStudy (solomonzelman):

apply the slope formula.... good luck!

OpenStudy (anonymous):

i try